OpenStudy (anonymous):

Find the limit as x approaches zero of (the square root of (x^6+9)-3)/x^6

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}\frac{ \sqrt{x^6+9}-3 }{ x^6 }\] please no l'hopital's rule. I'm suposed to be able to transform it algebraically. The answer is 1/6, I just don't know how to get there

OpenStudy (shubhamsrg):

Try to rationalize the numerator.

OpenStudy (anonymous):

what does that mean?

OpenStudy (shubhamsrg):

multiply and divide by sqrt(x^6 +9) + 3 See if it helps ?

OpenStudy (anonymous):

Ok i get to this point:

OpenStudy (anonymous):

\[\frac{ x^6+9-3 }{ x^6(\sqrt{x^6+9}+3) }\]

OpenStudy (anonymous):

Is that right and then how do i get past that?

OpenStudy (shubhamsrg):

In the numerator, in place of -3, there should be -9 , check that.

OpenStudy (anonymous):

got it! thanks!

OpenStudy (shubhamsrg):

Glad you did. ^_^